# Kernel-based Koopman approximants for control: Flexible sampling, error analysis, and stability

### v1

@@ 3.1. Inheritance of stability properties
Next, invoking the novel approximation bound of Theorem 2.4 and the respective error estimate on the data-driven surrogate dynamics Equation 12, we present our first main result, i.e., that a Lyapunov function (and, thus, asymptotic stability) of the dynamics Equation DS implies practical asymptotic stability w.r.t. Equation 12 and vice versa. Herein, the practical region $P$ can be rendered to be an arbitrary small neighbourhood of the equilibrium if the fill distance $h_{\mathcal{X}}$ is sufficiently small.
@@
Let $x^{*}\in\mathcal{X}$ be an equilibrium w.r.t. the dynamics Equation DS given by $F\in\mathcal{C}_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n})$ with $k\geq 1$ and, thus, also of the data-driven surrogate Equation 12 represented by $\widehat{F}$. [^3] See Proposition 3.1. Here, we assume that $\psi_{i}(x)=x_{i}$ for $i\in[1:n]$. Furthermore, let the function $V:\Omega\to\mathbb{R}_{\geq 0}$ admit a modulus of continuity $\omega_{V}\in\mathscr{K}$, and define the sublevel set $V^{-1}(c_{\Omega}):=\{x\in\Omega\ |\ V(x)\leq c_{\Omega}\}$, where $c_{\Omega}>0$ is chosen such that $V^{-1}(c_{\Omega})\subset\Omega$ is closed. Then, the following two statements hold:
@@
    Let $V$ be a Lyapunov function w.r.t. the dynamics Equation DS on $\Omega$ and assume [^4] This assumption can be ensured by decreasing the level $c_{\Omega}$ of the chosen sublevel set $V^{-1}(c_{\Omega})$. that the decrease condition Equation 15 holds for all $x\in V^{-1}(c_{\Omega})$, which renders this (bounded ) set forward-invariant w.r.t. Equation DS.
    Then, $x^{*}$ is practically asymptotically stable w.r.t. Equation 12 in the sense that, for every $\varepsilon>0$, the practical region $P$ can be chosen as a subset of the $\varepsilon$-ball $\mathcal{B}_{\varepsilon}(x^{*})$ if the fill distance $h_{\mathcal{X}}$ is sufficiently small.
@@
    The statement (i) holds upon switching the roles of $F$ and $\widehat{F}$, i.e., the existence of a Lyapunov function w.r.t. the data-driven surrogate dynamics Equation 12 implies practical asymptotic stability w.r.t. the dynamics Equation DS.
@@
We begin with assertion (i). Hence, let $V$ be a Lyapunov function w.r.t. the dynamics Equation DS. Since $\psi_{i}(x)=x_{i}$ for $i\in[1:n]$, it follows that $\Upsilon$ can be chosen such that $\omega_{\Upsilon}(r)=r$, cf. Remark 2.7. Then, for all $x\in Y:=V^{-1}(c_{\Omega})$, we have
@@
Equation 15superscriptEquation ${{\leq}}$ $\omega_{V}(\|\widehat{F}(x)-F(x)\|)+V(x)-\alpha_{V}(\|x-x^{*}\|)$
Corollary 2.8superscriptCorollary ${{\leq}}$ $\Big{[}\omega_{V}(\bar{C}h_{\mathcal{X}}^{k+1/2})-s\alpha_{V}(\|x-x^{*}\|)\Big{]}+V(x)-\underbrace{(1-s)\alpha_{V}(\|x\|)}_{:=\widehat{\alpha}(\|x-x^{*}\|)},$
where $\widehat{\alpha}\in\mathscr{K}$, $s\in(0,1)$, and $\bar{C}:=C\|\Psi\|_{\mathcal{N}^{M}_{\Phi_{n},k}(\Omega)}$ with $C>0$ from Corollary 2.8. Hence, if the inequality
@@
holds, we have a Lyapunov-decrease inequality of the form Equation 15 along the data-driven surrogate dynamics Equation 12.
Next, assume w.l.o.g. that $\varepsilon\in(0,\alpha_{2}^{-1}(c_{\Omega})]$ such that $B_{\varepsilon}(x^{*})\subset V^{-1}(c_{\Omega})$. Indeed, if $x\in\mathcal{B}_{\varepsilon}(x^{*})$, then $\|x-x^{*}\|\leq\alpha_{2}^{-1}(c_{\Omega})$ and, thus, $\alpha_{2}(\|x-x^{*}\|)\leq c_{\Omega}$. Consequently, $x$ is an element of $V^{-1}(c_{\Omega})$ in view of the second inequality in Equation 14.
We now construct a forward invariant set $P\subset B_{\varepsilon}(x^{*})$ such that Equation 17 holds on $Y\setminus P$. This then implies the claim due to Proposition 3.4. To this end, let $c_{\varepsilon}>0$ such that $P:=V^{-1}(c_{\varepsilon})\subset\mathcal{B}_{\varepsilon}(x^{*})$ holds and choose $\eta>0$ small enough to ensure the inclusion
$\widehat{F}(\mathcal{B}_{\eta}(x^{*})):=\{\widehat{F}(x):x\in\mathcal{B}_{\eta}(x^{*})\}\subset P$ (18)
using $\widehat{F}(x^{*})=x^{*}$, the continuity of the map $\widehat{F}$ and Condition (14) for the Lyapunov function $V$. Then, choose $c_{\eta}>0$ maximal such that the inclusion $V^{-1}(c_{\eta})\subset\mathcal{B}_{\eta}(x^{*})$ holds. By definition, for every $x\in V^{-1}(c_{\eta})$, we have $x\in\mathcal{B}_{\eta}(x^{*})$ and, thus, $\widehat{F}(x)\in V^{-1}(c_{\varepsilon})$, i.e., $V(\widehat{F}(x))\leq c_{\varepsilon}$. Thus, also $\widehat{F}(x)\in\mathcal{B}_{\varepsilon}(x^{*})$.
@@
Finally, for every $x\in V^{-1}([c_{\eta},c_{\Omega}])=\{x\in\mathbb{R}^{n}:c_{\eta}\leq V(x)\leq c_{\Omega}\}$, choosing the fill distance $h_{\mathcal{X}}$ sufficiently small [^5] Recall that $\omega_{V}\in\mathscr{K}$ is strictly increasing. Then, for instance, a fill distance $h_{\mathcal{X}}$ satisfying $h_{\mathcal{X}}^{k+1/2}\leq\omega_{V}^{-1}(s\alpha_{V}(\|\delta\|))\bar{C}^{-1}D^{-1}$ with $\delta>0$ such that $\mathcal{B}_{\delta}(x^{*})\subset V^{-1}(c_{\eta})$ and $D$ satisfying $V^{-1}(c_{\Omega})\subset\mathcal{B}_{D}(x^{*})$ suffices., we get forward invariance of the set $P$ and the required Lyapunov decrease on $S:=Y\setminus P$ showing that $x^{*}$ is practically asymptotically stable with the practical region $P$ contained in the desired $\varepsilon$-ball.
@@
To complete the proof, we show assertion (ii). If $V$ is a Lyapunov function for the data-driven surrogate dynamics Equation 12, the proof follows analogously as in the previous case (i) with $F$ instead of $\widehat{F}$. ∎
We conclude this subsection by providing two extensions of this result.
Remark 3.6.
@@
(a) In principle, we could have shown semi-global (practical) asymptotic stability. To this end, one needs a global Lyapunov function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}_{\geq 0}$. Then, for a given compact set $K$, we determine a sublevel set $Y$ large enough such that the inclusion $K\subset Y$ holds using that $V$ is proper due to Equation 14. Then, we set $\Omega$ such that $Y\subset\Omega$ and generate the respective data-driven surrogate model.
@@
(b) Moreover, the results can be directly transferred to the dynamics $\widehat{F}_{\lambda}$ defined in Equation 10, resulting from the regularized regression problem, that is, with $\lambda>0$. The only change in the proof is the choice of $\eta$ such that the inclusion Equation 18 holds. Here, the regularization parameter has to be chosen sufficiently small such that the equilibrium $x^{*}$ of the dynamics $F$ is almost preserved, i.e., such that the inclusion still holds for the desired precision $\varepsilon$.
@@ 3.2. Proportional bounds and asymptotic stability
In this part, we provide conditions under which asymptotic stability is preserved. To this end, we first present novel proportional bounds on the approximation error extending previous results from [4, 43], where comparable bounds were obtained under additional assumptions in order to cover the projection error and which only hold with a certain probability. To this end, we refine Theorem 2.4 to reflect the objective of stabilization by means of proportionality of the right-hand side when approaching the equilibrium. Then, assuming some compatibility condition linking the modulus of continuity $\omega_{V}$ and the Lyapunov decrease condition Equation 15, we show that asymptotic stability is preserved for the data-driven surrogate model and vice versa.
@@
with $\widehat{F}$ defined in Equation 12, where $\varepsilon>0$ may be rendered arbitrarily small by using more data points and $\omega_{\Upsilon}$ denotes the modulus of continuity of the left-inverse $\Upsilon$ of $\Psi$.
In view of the desired bound Equation 19, observe that the previously established estimate of Corollary 2.8 is sub-optimal in the case $\lambda=0$: For data points $x\in\mathcal{X}$, the left-hand side in the estimate Equation 11 vanishes while the right-hand side remains constant. Thus, the next theorem improves the error bounds of Corollary 2.8 by including a dependence on the distance to the data set in the upper bound. In its proof and in the sequel, we work with the spaces of continuous and bounded functions with continuous and bounded derivatives defined as follows. For $s\in\mathbb{N}$, we consider the space $C_{b}^{s}(\Omega)$ of continuous functions $f:\Omega\to\mathbb{R}$ for which $D^{\alpha}f$ is bounded on $\Omega$ for all multiindices $\alpha\in\mathbb{N}_{0}^{d}$, $|\alpha|\leq s$, endowed with the norm
$\|f\|_{C_{b}^{s}(\Omega)}:=\sum_{|\alpha|\leq s}\,\sup_{x\in\Omega}|D^{\alpha}f(x)|=\sum_{|\alpha|\leq s}\|D^{\alpha}f\|_{C_{b}(\Omega)}.$
@@
Further, if additionally $F\in C_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n})$ holds for the dynamics Equation DS, there are constants $C,h_{0}>0$ such that, for any finite set $\mathcal{X}=\{x_{i}\}_{i=1}^{d}\subset\Omega$ of sample points with $h_{\mathcal{X}}\leq h_{0}$, we have
@@
We begin with the proof of the first inequality w.r.t. interpolation errors. For given $f\in\mathcal{N}_{\Phi_{n,k}}(\Omega)$, we define the error function by $e_{f}(x):=f(x)-(P_{\mathcal{X}}f)(x)$, $x\in\Omega$. Since $k\geq 1$, it follows from Sobolev’s embedding theorem (see, e.g., [1, Theorem 4.12]) that $\mathcal{N}_{\Phi_{n,k}}(\Omega)=H^{\sigma_{n,k}}(\Omega)\subset C_{b}^{1}(\Omega)$ and hence $e_{f}\in C_{b}^{1}(\Omega)$. For arbitrary $j\in[1:d]$, we infer
$|e_{f}(x)|=|e_{f}(x)-e_{f}(x_{j})|\leq\Big{(}\sup_{z\in\Omega}\,\|\nabla e_{f}(z)\|\Big{)}\cdot\|x-x_{j}\|.$
As this holds for all $j\in[1:d]$, we conclude $|e_{f}(x)|\leq(\sup_{z\in\Omega}\,\|\nabla e_{f}(z)\|)\operatorname{dist}(x,\mathcal{X})$. Now, from Lemma 2.5 with $\lambda=0$ and $|\alpha|=1$ corresponding to all first derivatives, we obtain
$\|\nabla e_{f}(z)\|^{2}=\sum_{i=1}^{n}|\partial_{i}e_{f}(z)|^{2}\,\leq\,nC^{2}h_{\mathcal{X}}^{2k-1}\|f\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)}^{2}$
for all $z\in\Omega$. The claim follows after redefining $C:=\sqrt{n}C$.
Next, we show the second claim. Due to [20, Theorem 4.2], $F\in C_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n})$ implies the invariance of the native space $\mathcal{N}_{\Phi_{n,k}}(\Omega)$ under the Koopman operator $\mathcal{K}$, and the restricted operator $\mathcal{K}:\mathcal{N}_{\Phi_{n,k}}(\Omega)\to\mathcal{N}_{\Phi_{n,k}}(\Omega)$ is bounded. Therefore, for $f\in\mathcal{N}_{\Phi_{n,k}}(\Omega)$ and $x\in\Omega$, we may use the just established result to estimate
@@
$\leq C\|\mathcal{K}\|_{\mathcal{N}_{\Phi_{n,k}}\to\mathcal{N}_{\Phi_{n,k}}}h_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})\|f\|_{\mathcal{N}_{\Phi_{n,k}}},$
which proves the claim. ∎
We now provide a proportional error bound for the surrogate dynamics Equation 12 as claimed in Equation 19.
@@
Under the assumptions of Theorem 3.7, there are constants $C,h_{0}>0$ such that for any finite set $\mathcal{X}=\{x_{i}\}_{i=1}^{d}\subset\Omega$ of sample points with $h_{\mathcal{X}}\leq h_{0}$ we have
$\|F(x)-\widehat{F}(x)\|\leq\omega_{\Upsilon}\big{(}Ch_{\mathcal{X}}^{k-1/2}\|\Psi\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)^{M}}\operatorname{dist}(x,\mathcal{X})\big{)}.$
In particular, for any $x^{*}\in\mathcal{X}$,
$\|F(x)-\widehat{F}(x)\|\leq\omega_{\Upsilon}\big{(}Ch_{\mathcal{X}}^{k-1/2}\|\Psi\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)^{M}}\|x-x^{*}\|\big{)}.$
@@
The first claim follows from the proof of Corollary 2.8 using Theorem 3.7 in the second-last estimate. The second claim follows directly from the definition of the distance. ∎

### v3

@@ 3.1. Inheritance of stability properties
Next, using the novel approximation bound of theorem 2.4 and the error estimate on the surrogate dynamics eq. 11, we present our first main result, i.e., that a Lyapunov function (and, thus, asymptotic stability) of the dynamics eq. DS implies practical asymptotic stability w.r.t. eq. 11 and vice versa. Herein, the practical region $P$ can be rendered to be an arbitrary small neighbourhood of the equilibrium if the fill distance $h_{\mathcal{X}}$ is sufficiently small.
@@
Let $x^{*}\in\mathcal{X}$ be an equilibrium w.r.t. the dynamics eq. DS given by $F\in\mathcal{C}_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n})$ with $\sigma_{n,k}$ defined by eq. 3 and $k\geq 1$ and, thus, also of the data-driven surrogate eq. 11 represented by $\widehat{F}$ (see proposition 3.1). Furthermore, let $\psi_{i}(x)=x_{i}$, $i\in[1:n]$, be given. Then, the following two statements hold:
@@
    Let $V:\Omega\to\mathbb{R}_{\geq 0}$ be a Lyapunov function w.r.t. the dynamics eq. DS on $\Omega$ with modulus of continuity $\omega_{V}\in\mathscr{K}$. Further, let $c_{\Omega}>0$ be sufficiently small such that the sublevel set $V^{-1}(c_{\Omega}):=\{x\in\Omega\mid V(x)\leq c_{\Omega}\}$ is closed and the decrease condition eq. 15 holds for all $x\in V^{-1}(c_{\Omega})$. [^3] In particular, $V^{-1}(c_{\Omega})$ is, then, forward invariant w.r.t. eq. DS. Then, $x^{*}$ is practically asymptotically stable w.r.t. eq. 11 in the sense that, for every $\varepsilon\in(0,\alpha_{2}^{-1}(c_{\Omega})]$, the practical region $P$ can be chosen as a subset of the $\varepsilon$-ball $\mathcal{B}_{\varepsilon}(x^{*})$ if only the fill distance $h_{\mathcal{X}}$ is sufficiently small.
@@
    The statement (i) holds upon switching the roles of $F$ and $\widehat{F}$, if the comparison functions of the Lyapunov function are uniform for all $h_{\mathcal{X}}\in(0,h_{0}]$ for some $h_{0}$ [^4] This assumption holds, e.g., for value functions in data-driven predictive control, see [5]. The proof follows very similarly to the proof of lemma 3.6 using the uniform error bound., i.e., the existence of a Lyapunov function w.r.t. the data-driven surrogate dynamics eq. 11 uniform in the fill distance implies practical asymptotic stability w.r.t. the dynamics eq. DS.
The imposed smallness condition on the fill distance $h_{\mathcal{X}}$ corresponds to a condition on the collected data, which may be easily met by grid-based sampling in the set $\Omega$ as illustrated in our numerical simulations in Section 3.3, but is typically not satisfied when sampling asymptotically stable systems along trajectories.
To prove theorem 3.5, we require Lipschitz continuity of the surrogate model, uniformly w.r.t. the approximation accuracy depending on the fill distance $h_{\mathcal{X}}$ and the regularization parameter $\lambda$. To outline a possible extension of theorem 3.5 to the approximant $F_{\lambda}$ generated with regularization parameter $\lambda>0$, we show the respective assertion for this more general case. Hereby, it becomes apparent that the regularization parameter $\lambda$ has to be chosen proportional to the squared fill distance $h_{\mathcal{X}}$.
Lemma 3.6 (Lipschitz continuity of the data-driven surrogate $\widehat{F}_{\lambda}$).
Let $k\geq 1$ and $F\in C_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n})$ with $\sigma_{n,k}:=\frac{n+1}{2}+k$. Then, there exist constants $h_{0}>0,\bar{L}\geq 0$ such that, if the fill distance $h_{\mathcal{X}}$ and the regularization parameter $\lambda$ satisfy $h_{\mathcal{X}}\leq h_{0}$ and $h_{\mathcal{X}}^{-1}\sqrt{\lambda}\leq c_{\lambda}$ for some $c_{\lambda}\in[0,\infty)$, the surrogate $\widehat{F}_{\lambda}$ defined by (11) is Lipschitz continuous in $x$ with Lipschitz constant $L_{\widehat{F}_{\lambda}}\leq\bar{L}$ , i.e.,
$\|\widehat{F}_{\lambda}(x)-\widehat{F}_{\lambda}(y)\|\leq L_{\widehat{F}_{\lambda}}\|x-y\|\qquad\forall\,x,y\in\Omega.$ (16)
@@
$\widehat{F}_{\lambda}(x)-\widehat{F}_{\lambda}(y)=[\widehat{F}_{\lambda}(y)+(x-y)^{\top}J_{\widehat{F}_{\lambda}}(\xi)]-\widehat{F}_{\lambda}(y)=(x-y)^{\top}J_{\widehat{F}_{\lambda}}(\xi)$ holds for $x,y\in\Omega$ and some $\xi\in\{(t-1)x+ty:t\in[0,1]\}$ by Taylor’s theorem, where $J$ stands for the Jacobian. Hence, we may estimate
$\|\widehat{F}_{\lambda}(x)\hskip-1.42262pt-\hskip-1.42262pt\widehat{F}_{\lambda}(y)\|\leq\|x\hskip-1.42262pt-\hskip-1.42262pty\|\|J_{\widehat{F}_{\lambda}}(\xi)\|\leq\|x\hskip-1.42262pt-\hskip-1.42262pty\|\Big[\|J_{\widehat{F}_{\lambda}}(\xi)\hskip-1.42262pt-\hskip-1.42262ptJ_{F}(\xi)\|+\|J_{F}(\xi)\|\Big].$
Since the function $F$ is continuously differentiable on the compact set $\Omega$, we have $\|J_{F}(\xi)\|\leq\widetilde{C}\in(0,\infty)$. To compute a bound for $\|J_{\widehat{F}_{\lambda}}(\xi)-J_{F}(\xi)\|$, we apply lemma 2.5 with multiindex $|\alpha|=1$ to obtain $\|J_{\widehat{F}_{\lambda}}(z)-J_{F}(z)\|\leq C_{J_{F}}(h_{\mathcal{X}}^{k-1/2}+h_{\mathcal{X}}^{-1}\sqrt{\lambda})$, where $C_{J_{F}}$ is a constant independent of the fill distance $h_{\mathcal{X}}$ and $\lambda$. Overall, we get the desired inequality (16) with $L_{\widehat{F}}$ given by $\widetilde{C}+C_{J_{F}}(h_{\mathcal{X}}^{k-1/2}+h_{\mathcal{X}}^{-1}\sqrt{\lambda})$. Hence, the assertion follows with $\bar{L}:=\widetilde{C}+C_{J_{F}}(h_{0}^{k-1/2}+c_{\lambda})$. ∎
The continuity in lemma 3.6 is uniform in the approximation bound $\varepsilon>0$. This is in contrast to finite-element dictionaries, see [49], where the derivative of the ansatz functions increases for decreasing mesh size. Here, an error bound on the derivative for kernel-based interpolations is central for proving uniform estimates. Moreover, we note that the adaptive choice of the regularization parameter (depending on the fill distance) stated in Lemma 3.6 is closely related to the approach of [56].
Next, using Lemma 3.6, we prove Theorem 3.5.
Proof.
We begin with assertion (i) meaning that $V$ is a Lyapunov function w.r.t. the dynamics eq. DS. Since $\psi_{i}(x)=x_{i}$ for all $i\in[1:n]$, it follows that $\Upsilon$ can be chosen such that $\omega_{\Upsilon}(r)=r$, cf. remark 2.7 and the subsequent comment. Then, for all $x\in Y:=V^{-1}(c_{\Omega})$, we have using eq. 15 and corollary 2.8 constant $C>0$
@@
$\leq$ $\omega_{V}(\|\widehat{F}(x)-F(x)\|)+V(x)-\alpha_{V}(\|x-x^{*}\|)$
$\leq$ $\Big[\omega_{V}(\bar{C}h_{\mathcal{X}}^{k+1/2})-s\alpha_{V}(\|x-x^{*}\|)\Big]+V(x)-(1-s)\alpha_{V}(\|x{-x^{*}}\|)$
where $s\in(0,1)$, $(1-s)\alpha_{V}\in\mathscr{K}$, and $\bar{C}:=C\|\Psi\|_{\mathcal{N}^{M}_{{\Phi_{n,k}}}(\Omega)}$. Hence, if
@@
holds, we have a Lyapunov-decrease inequality of the form eq. 15 along the data-driven surrogate dynamics eq. 11. Before we proceed, we point out that the inclusion $B_{\varepsilon}(x^{*})\subset V^{-1}(c_{\Omega})$ holds due to our choice $\varepsilon\in(0,\alpha_{2}^{-1}(c_{\Omega})]$: If $x\in\mathcal{B}_{\varepsilon}(x^{*})$, then $\|x-x^{*}\|\leq\alpha_{2}^{-1}(c_{\Omega})$ and, thus, $\alpha_{2}(\|x-x^{*}\|)\leq c_{\Omega}$. Consequently, $x$ is an element of $V^{-1}(c_{\Omega})$ in view of the second inequality in eq. 14.
@@
We construct a forward invariant set $P\subset B_{\varepsilon}(x^{*})$ such that eq. 18 holds on $Y\setminus P$. This implies the claim due to proposition 3.4. To this end, let $c_{\varepsilon}>0$ be such that $P:=V^{-1}(c_{\varepsilon})\subset\mathcal{B}_{\varepsilon}(x^{*})$ holds and choose $\eta>0$ small enough to ensure the inclusion $\widehat{F}(\mathcal{B}_{\eta}(x^{*})):=\{\widehat{F}(x):x\in\mathcal{B}_{\eta}(x^{*})\}\subset P$ using $\widehat{F}(x^{*})=x^{*}$, the uniform Lipschitz continuity of the map $\widehat{F}$ shown in lemma 3.6, and Condition (14) for the Lyapunov function $V$. Then, choose $c_{\eta}>0$ maximal such that the inclusion $V^{-1}(c_{\eta})\subset\mathcal{B}_{\eta}(x^{*})$ holds. By definition, for every $x\in V^{-1}(c_{\eta})$, we have $x\in\mathcal{B}_{\eta}(x^{*})$ and, thus, $\widehat{F}(x)\in V^{-1}(c_{\varepsilon})$, i.e., $V(\widehat{F}(x))\leq c_{\varepsilon}$. Thus, also $\widehat{F}(x)\in\mathcal{B}_{\varepsilon}(x^{*})$. Finally, for every $x\in V^{-1}([c_{\eta},c_{\Omega}])=\{x\in\mathbb{R}^{n}:c_{\eta}\leq V(x)\leq c_{\Omega}\}$, choosing the fill distance $h_{\mathcal{X}}$ sufficiently small [^5] Recall that $\omega_{V}\in\mathscr{K}$ is strictly increasing. Then, for instance, a fill distance $h_{\mathcal{X}}$ satisfying $h_{\mathcal{X}}^{k+1/2}\leq\omega_{V}^{-1}(s\alpha_{V}(\|\delta\|))\bar{C}^{-1}D^{-1}$ with $\delta>0$ such that $\mathcal{B}_{\delta}(x^{*})\subset V^{-1}(c_{\eta})$ and $D$ satisfying $V^{-1}(c_{\Omega})\subset\mathcal{B}_{D}(x^{*})$ suffices., we get forward invariance of the set $P$ and the required Lyapunov decrease on $S:=Y\setminus P$ showing that $x^{*}$ is practically asymptotically stable with the practical region $P$ contained in the desired $\varepsilon$-ball.
@@
Next, we show assertion (ii). If $V$ is a Lyapunov function for the data-driven surrogate dynamics eq. 11, the proof follows analogously as in case (i) with $F$ instead of $\widehat{F}$. The uniformity of the comparison functions ensures that the involved quantities do not deteriorate for small fill distances, e.g., that (18) is satisfied in any neighborhood of the equilibrium when choosing the fill distance small enough. ∎
@@
We conclude this subsection by providing two extensions of this result. First, we could have shown semi-global (practical) asymptotic stability: To this end, one needs a global Lyapunov function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}_{\geq 0}$. Then, for a given compact set $K$, we determine a sublevel set $Y$ large enough such that the inclusion $K\subset Y$ holds using that $V$ is proper due to eq. 14. Then, we set $\Omega$ such that $Y\subset\Omega$ and generate the respective data-driven surrogate model.
@@
Second, the results can be transferred to the dynamics $\widehat{F}_{\lambda}$ defined in eq. 9, resulting from the regularized regression problem ($\lambda>0$). The only change in the proof is the choice of $\eta$ to ensure $\widehat{F}(\mathcal{B}_{\eta}(x^{*}))\subset P$. Here, the regularization parameter has to be chosen sufficiently small such that this inclusion is preserved for the given $\varepsilon$.
@@ 3.2. Proportional bounds and asymptotic stability
We provide conditions under which asymptotic stability is preserved. First, we present novel proportional bounds on the approximation error extending previous results from [4, 52], where comparable, yet probabilistic bounds were obtained under additional assumptions to cover the projection error. To this end, we refine theorem 2.4 to reflect the objective of stabilization by means of proportionality of the right-hand side when approaching the equilibrium. Then, assuming some compatibility condition linking the modulus of continuity $\omega_{V}$ and the Lyapunov decrease condition eq. 15, we show that asymptotic stability is preserved for the data-driven surrogate model and vice versa.
@@
with $\widehat{F}$ defined in eq. 11, where $\varepsilon>0$ may be rendered arbitrarily small by using more data points and $\omega_{\Upsilon}$ denotes the modulus of continuity of the left-inverse $\Upsilon$ of $\Psi$.
In view of the desired bound eq. 19, observe that the previously established estimate of corollary 2.8 is sub-optimal in the case $\lambda=0$: For data points $x\in\mathcal{X}$, the left-hand side in the estimate eq. 10 vanishes while the right-hand side remains constant. Thus, the next theorem improves the error bounds of corollary 2.8 by including a dependence on the distance to the data set in the upper bound. In its proof and in the sequel, we work with the spaces of continuous and bounded functions with continuous and bounded derivatives defined as follows. For $s\in\mathbb{N}$, we consider the space $C_{b}^{s}(\Omega)$ of continuous functions $f:\Omega\to\mathbb{R}$ for which $D^{\alpha}f$ is bounded on $\Omega$ for all multiindices $\alpha\in\mathbb{N}_{0}^{d}$, $|\alpha|\leq s$, endowed with the norm $\|f\|_{C_{b}^{s}(\Omega)}:=\sum\nolimits_{|\alpha|\leq s}\,\sup\nolimits_{x\in\Omega}|D^{\alpha}f(x)|=\sum\nolimits_{|\alpha|\leq s}\|D^{\alpha}f\|_{C_{b}(\Omega)}.$
@@
Further, if additionally $F\in C_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n})$ for eq. DS, there are constants $C,h_{0}>0$ such that, for any finite set $\mathcal{X}=\{x_{i}\}_{i=1}^{d}\subset\Omega$ of sample points with $h_{\mathcal{X}}\leq h_{0}$, we have
@@
First, w prove the first inequality w.r.t. interpolation errors. For given $f\in\mathcal{N}_{\Phi_{n,k}}(\Omega)$, we define the error function by $e_{f}(x):=f(x)-(P_{\mathcal{X}}f)(x)$, $x\in\Omega$. Since $k\geq 1$, it follows from Sobolev’s embedding theorem (see, e.g., [1, Theorem 4.12]) that $\mathcal{N}_{\Phi_{n,k}}(\Omega)=H^{\sigma_{n,k}}(\Omega)\subset C_{b}^{1}(\Omega)$ and hence $e_{f}\in C_{b}^{1}(\Omega)$. Using $e_{f}(x_{j})=0$, we infer $|e_{f}(x)|=|e_{f}(x)-e_{f}(x_{j})|\leq(\sup_{z\in\Omega}\,\|\nabla e_{f}(z)\|)\|x-x_{j}\|$ for arbitrary $j\in[1:d]$. As this holds for all $j\in[1:d]$, we conclude $|e_{f}(x)|\leq(\sup_{z\in\Omega}\,\|\nabla e_{f}(z)\|)\operatorname{dist}(x,\mathcal{X})$. Now, from lemma 2.5 with $\lambda=0$ and $|\alpha|=1$ corresponding to all first derivatives, we obtain $\|\nabla e_{f}(z)\|^{2}=\sum_{i=1}^{n}|\partial_{i}e_{f}(z)|^{2}\,\leq\,nC^{2}h_{\mathcal{X}}^{2k-1}\|f\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)}^{2}$ for all $z\in\Omega$. The claim follows after redefining $C:=\sqrt{n}C$.
Next, we show the second claim. Due to [28, Theorem 4.2], $F\in C_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n})$ implies the invariance of the native space $\mathcal{N}_{\Phi_{n,k}}(\Omega)$ under the Koopman operator $\mathcal{K}$, and the restricted operator $\mathcal{K}:\mathcal{N}_{\Phi_{n,k}}(\Omega)\to\mathcal{N}_{\Phi_{n,k}}(\Omega)$ is bounded. Therefore, for $f\in\mathcal{N}_{\Phi_{n,k}}(\Omega)$ and $x\in\Omega$, we use the just established result to estimate to prove the claim:
@@
$\leq C\|\mathcal{K}\|_{\mathcal{N}_{\Phi_{n,k}}\to\mathcal{N}_{\Phi_{n,k}}}h_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})\|f\|_{\mathcal{N}_{\Phi_{n,k}}}.$
We now provide a proportional error bound for the surrogate dynamics eq. 11 as claimed in eq. 19.
@@
Under the assumptions of theorem 3.7, there are constants $C,h_{0}>0$ such that for any finite set $\mathcal{X}=\{x_{i}\}_{i=1}^{d}\subset\Omega$ of sample points with $h_{\mathcal{X}}\leq h_{0}$ we have $\|F(x)-\widehat{F}(x)\|\leq\omega_{\Upsilon}\big(Ch_{\mathcal{X}}^{k-1/2}\|\Psi\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)^{M}}\operatorname{dist}(x,\mathcal{X})\big)$ and, in particular,
$\|F(x)-\widehat{F}(x)\|\leq\omega_{\Upsilon}\big(Ch_{\mathcal{X}}^{k-1/2}\|\Psi\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)^{M}}\|x-x^{*}\|\big)\qquad\forall\,x^{*}\in\mathcal{X}.$
@@
The first claim follows from the proof of corollary 2.8 using theorem 3.7 in the second-last estimate, while the second directly from the definition of the distance. ∎